OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x2 - 3 and g(x) = square root of 3+x

OpenStudy (anonymous):

please help!!!!!!! im on a timed test please

OpenStudy (precal):

sub one function into the other and simplify then reverse your process

OpenStudy (anonymous):

i just can seem to get it to work with my cal:/

OpenStudy (anonymous):

im terrible at math!

OpenStudy (precal):

you will be using your algebra skills to simplify

Parth (parthkohli):

Test: what is \(\rm f(g(x))\) if \(\rm f(x) = x+1\) and \(\rm g(x) = x+2\)?

OpenStudy (anonymous):

f(x+2)=x+3 ??

OpenStudy (precal):